40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Message ID: 2406     Entry time: Mon Aug 19 13:57:02 2019
Author: Shalika 
Type: Noise Budget 
Category: PD noise 
Subject: Dark noise of Sb3513_A2 

[Koji, Nathan, Duo, Shalika]

The dark noise of the Sb3513_A2 photodiode was observed using an SRS785 spectrum analyzer. Different bias voltages were provided to understand the dependency of dark noise on the bias voltage. It was observed that as the bias voltage deeps on decreasing the dark noise decrease too. A transimpedance amplifier with a gain of 5k was used to convert photocurrent into voltage. A Sallen key low pass filter was used in order to provide a low noise bias. 

** Since the gain of the TIA was 5k so the output voltage noise was divided by 5k in order to get dark current noise.

Attachment 1: Dark current noise across Sb3513_A2 500um

Attachment 2: Dark current noise across Sb3513_A2 750um

Attachment 3: Dark current noise across Sb3513_A2 1000um

Attachment 4: Dark Current across Sb3513_A2 500um, 750um, 1000um respectively

Attachment 5: zip file containing all data

[added (Nathan)] Attachment 6: Experiment setup.

Attachment 1: Noise_across_500um.pdf  140 kB  Uploaded Mon Aug 26 11:52:53 2019  | Hide | Hide all
Noise_across_500um.pdf
Attachment 2: Noise_across_750um.pdf  117 kB  Uploaded Mon Aug 26 11:53:07 2019  | Hide | Hide all
Noise_across_750um.pdf
Attachment 3: Noise_across_1000um.pdf  116 kB  Uploaded Mon Aug 26 11:53:18 2019  | Hide | Hide all
Noise_across_1000um.pdf
Attachment 4: Dark_current.pdf  24 kB  Uploaded Mon Aug 26 11:53:31 2019  | Hide | Hide all
Dark_current.pdf
Attachment 5: JPL_Sb35313_A2.zip  2.787 MB  Uploaded Mon Aug 26 11:53:44 2019
Attachment 6: PD_test_setup.pdf  25 kB  Uploaded Mon Aug 26 11:53:56 2019  | Hide | Hide all
PD_test_setup.pdf
ELOG V3.1.3-